# EEG Foundation Models: A Critical Review of Current Progress and Future Directions

### v2

@@ 4.1 Preparation and Representation of Input Data
SD-71 [95] (Sleep deprivation detection - binary) $eval^{\#}$
ADHD-Adult [96] (ADHD classification - binary) $eval^{\#}$
ADHD-Child [97] (ADHD classification - binary) $eval^{\#}$
Schizophrenia-28 [98] (Schizophrenia classification - binary) $eval^{\#}$
Depression-122 [99] (Depression classification - binary) $eval^{\#}$
MDD-64 [100] (Major depression detection - binary) $eval^{\#}$
AD-65 [101] (Alzheimer’s disease classification - binary) $eval^{\#}$
BCI Competition IV-1 [102] (Movement classification - binary) $train$ $train$
Emobrain [103] (Emotion classification - multiclass) $train$ $train$
Grasp/Lift EEG Challenge [104] (Grasp classification/regression) $train$ $train$
Inria BCI Challenge [105] (Next letter classification - multiclass) $train$ $train$
EEG Motor Imagery Dataset [106] (Classification - multiclass) $train$ $train$ $train$
Visual Categorization EEG Data [107] (Classification - multiclass) $train$ $train$
Resting State EEG Data  [108] (Eye state classification - binary) $train$ $train$
SEED Series [109, 110, 111] (Emotion classification - multiclass) $train/eval^{*}$ $train/eval^{*}$ $train/eval^{*}$
SPIS Resting State Dataset [112] (CVS and HRT regression) $train$ $train$
Brain-Invaders [113] (Target vs. non-target classification - binary) $train$ $train$
Multiple data sources [114, 115, 116, 117, 118] (Unknown tasks) $train$ $train$
MoBI [119] (Gait angle regression) $eval^{\#}$
BCI Competition IV Dataset 2a [120] $eval^{\#}$
Workload [121] (Workload classification - binary) $eval^{\#}$
@@
Brain TreeBank [80] (EEG classification - multiclass) $train/eval^{*}$
@@
Private dataset collection without source (Unknown) $train$
MAYO [122] (Seizure detection and multi-class classification) $eval^{\#}$ $train/eval^{*}$ $eval^{\#}$
FNUSA [122] (Seizure detection and multi-class classification) $eval^{\#}$ $train/eval^{*}$ $eval^{\#}$
CCEP [81] (Seizure onset zone localization - binary) $train$
@@ 4.2 Modeling and Pretraining
Architectural components: The model architecture and training objectives play a central role in determining the effectiveness of a foundation model in extracting information from the input data [8]. EEG-FMs utilized various components at different stages of the pipeline, including tokenizers, local representation learning modules, spatial and/or temporal attention modules to learn dependencies, and task heads. Almost all the EEG-FMs we review in this article are transformer-based, except Mentality, which uses a combination of convolutional layers and a mamba-based architecture. In the transformer-based models, the input signal is divided into a fixed number of small patches/tokens to construct the input sequence. Some models that utilized raw time series inputs leveraged convolutional layers to learn morphological features (e.g., Neuro-GPT, LaBraM, Mentality, NeuroLM).
Neural tokenizers: Some models, such as LaBraM, NeuroLM, and EEGFormer, integrated separate tokenizers to transform the input signals into discrete tokens or codes. These EEG-FMs adapted a VQ-VAE[123]-like architecture to train a neural tokenizer, which mapped input patches into a fixed set of discrete embeddings called the codebook. During tokenizer training, latent codes are mapped to nearest neighbor embeddings from the codebook to reconstruct some known feature of the input, such as power amplitudes and phase in LaBraM and temporal and spectral reconstruction in NeuroLM. All of these models utilized neural codebooks in the context of reconstructing raw time series inputs using a transformer-like architecture. While NeuroLM and EEGFormer utilized codebook embeddings to map the encoder outputs to discrete codes prior to decoding, LaBraM utilized pretrained codebook embeddings of input signal patches as targets during model pretraining.
Spatial encoding: EEG-FMs implemented different spatial encoding strategies, although they differ in how explicitly they handle spatial information. Brant and FoME, for example, used transformer encoders with spatial attention but without dedicated spatial embeddings. In contrast, LaBraM, BIOT, and NeuroLM injected spatial information through explicit spatial embeddings to enhance positional encoding and capture the topographic layout of EEG channels. Among them, only BIOT supported multiple electrode configurations, while LaBraM and NeuroLM are limited to a fixed 10–20 montage format for spatial encoding. Models such as Neuro-GPT and Mentality learned spatial dependencies implicitly using convolutional layers, without dedicated spatial embeddings. BrainWave incorporated spatial attention mechanisms to model inter-channel dependencies directly. In contrast, BrainBERT and EEGFormer omit spatial modeling altogether, using only transformer layers without convolution or spatial priors.
Self-supervision: SSL approaches are the key components that enable the development of foundation models by learning common features represented in massive unlabeled datasets[8]. Most EEG-FMs (seven out of ten) employed reconstruction of masked data as the primary SSL approach. EEGFormer model was pretrained by minimizing the reconstruction loss between input signals and signals decoded from codebook mappings of transformer-encoded input patches. NeuroLM employed an autoregressive approach to predict future patches from past patches, including EEG-text an alignment objective, and BIOT adopted a contrastive learning approach where the embeddings of masked inputs and the same inputs with different augmentations are minimized (i.e., the SimCLR [48] pretraining approach). Only one model – Neuro-GPT, employed masked data reconstruction in the latent space. Additionally, the LaBraM model also employed a masked-signal reconstruction approach; however, the codebook embeddings of the masked input patches from a previously-trained neural tokenizer were used as reconstruction targets.
Masking: The granularity of the learned representations in the EEG-FMs is influenced by the type and extent of masking used during self-supervised pretraining [124]. In models based on masked-signal reconstruction, masking is typically applied to random patches across temporal and spatial dimensions, as opposed to masking entire channels or full time segments. For example, Brant and FoME applied a $\sim$40uniform random masking, zeroing-out the masked input signal patches across time and channel dimensions. LaBraM similarly utilized a 50% patch-level masking, while BrainWave, despite operating in the spectrogram domain, performed masking entire spectrograms of several channels. On the other hand, BrainBERT applied a more fine-grained masking strategy, masking 5% from the time domain and 5% from the frequency domain, effectively masking $\sim$10of the input spectrogram. Mentality applied random channel masking, targeting specific channels rather than patches. NeuroLM introduced a stair-step masking scheme, in which next token is auto-regressively predicted in a channel-conditioned manner, whereas Neuro-GPT utilized a causal masking approach in the latent space, consistent with autoregressive generation.
@@
Patching and context length: The temporal resolution of tokens or patches, i.e., how much EEG data each patch contains, varies between models, reflecting different design choices in model architecture and target tasks. Brant and FoME utilized patches corresponding to six seconds of EEG, while BrainBERT utilized 5-second STFT/Superlet windows. Models such as LaBraM, NeuroLM, BIOT, and BrainWave adopted a finer resolution, using 1-second segments per patch, whereas Neuro-GPT utilized a patch size of two seconds. Similarly, context length also varied among EEG-FMs. FoME was trained using 90-second segments represented by 15 tokens. LaBraM could take up to 256 tokens as input, with the number of temporal tokens depending on the number of channels used. Neuro-GPT utilized a 32-token context window with overlapping segments, covering $\sim$57.8 seconds of EEG. BIOT utilized 19 tokens per channel and EEGFormer utilized 12 seconds of EEG context as input.
@@
Model scale: The number of trainable parameters (or weights) varied significantly between the ten EEG-FMs, ranging from 3.3M in BIOT to 1.7B in NeuroLM, although most models fall within the 200-500M parameter range. Some EEG-FMs, such as FoME, LaBraM, and NeuroLM, were developed in multiple scales with the same architecture, allowing users to choose a model scale according to their specific resource constraints.
@@ 4.3 Model Evaluations
Downstream task evaluations: Performance on downstream tasks post-adaptation was the primary metric used for evaluation in all the EEG-FMs. However, the clinical and non-clinical downstream tasks used for evaluation and the adaptation techniques varied significantly between them. The various datasets used for evaluation are shown in Table 2. While most EEG-FM evaluations were restricted to either scalp EEG or intracranial EEG, two models, namely FoME and BrainWave, performed evaluations on both modalities. Furthermore, most FMs utilized common adaptation techniques, such as linear probing and fine-tuning, while some FMs utilized more esoteric approaches, such as proto-type-based few-shot evaluations and instruction tuning.
@@
Almost all EEG-FMs, except Mentality, investigated fine-tuning for downstream task evaluations, where the decoder is generally discarded and the encoder is retained for additional training. On the other hand, five EEG-FMs (i.e., Brant, BrainBERT, LaBraM, EEGFormer, and Neuro-GPT) evaluated linear probing, where the encoder is typically frozen and additional task-specific heads on top of the encoder are trained using labeled downstream data. The BrainWave model was evaluated using a prototype-based few shot learning, where class prototypes were created by averaging the latent representations of a few examples from each class and downstream classifications were performed by measuring the similarities to those prototypes in the learned latent space. Furthermore, NeuroLM, which was pretrained using an EEG-text alignment objective, was evaluated using a text prompting approach. The model was fine-tuned (instruction-tuned) using an auto-regressive prediction task utilizing EEG-text combined tokens, where the text tokens contained both class prompts and labels. Subsequently, the model was evaluated on various tasks by providing queries containing EEG tokens and prompts, for which the model predicted the class labels. In the following, we briefly summarize the evaluations performed by each EEG-FM and the results.
@@
BrainBERT: This model was trained and evaluated on the same intracranial EEG dataset, Brain TreeBank. The downstream task focused on predicting the features of the movies subjects were watching during the EEG recording. Fine-tuning the pretrained BrainBERT model on this task substantially outperformed fully-supervised linear and multi-layer neural network models. Linear probing also achieved Area Under the Receiver Operating Characteristic Curve (AUC) values similar to the fully-supervised models.
@@
Neuro-GPT: This model was evaluated on a single scalp EEG dataset that was not used during pretraining, where the downstream task was a 4-class motor-imagery classification. The fully-supervised baselines considered in this study include BENDR [125], SVM [126], EEGNet [74], CTCNN [127], CCNN [128], and NG-CRAM [129]. When used in a supervised setting without any pretraining, this model performed similar to other supervised baselines. However, when the pretrained model was fine-tuned for the downstream task, it provided significant improvements in certain configurations. In contrast, linear probing performed worse than fully-supervised baselines.
@@
Brant: This model was evaluated on three intracranial EEG datasets, one of which was used for pretraining. The downstream tasks evaluated include signal forecasting, imputation, and seizure detection. For all these tasks, the pretrained model was fine-tuned or linear probed for the respective task. The fully-supervised baselines with pretraining considered in this study included (RP, TS, CPC) [7], BENDR [125], MVTS [130] and BrainBERT that are designed for brain signals and CoST [131], TF-C [132], PatchTST [133], TS-TCC [134] that are for general time series. Fully-supervised baselines without any pretraining included, handcrafted features-based methods such as spectral power [135], rythmicity spectrogram [136], and amplitude-integrated EEG [137] along with SEEG-Net [138]. In signal forecasting and imputation tasks, fine-tuned Brant consistently outperformed several time series baselines and its own linear probing version. In the seizure detection task, both the fine-tuned and linear probed versions of Brant outperformed other supervised and EEG-FM baselines, including BrainBERT. These observations held true even in low-labeled settings.
@@
BIOT: This model was evaluated entirely on scalp EEG datasets on various clinical tasks. The supervised baselines considered in this study included SPaRCNet [139], ContraWR [140], CNN-Transformer [141], FFCL [142], ST-Transformer [143]. Evaluations showed that the pretrained BIOT model with fine-tuning outperformed several fully-supervised baselines in CHB-MIT seizure detection and TUAB normal/abnormal classification. Whereas, in two other tasks, namely IIIC seizure classification and TUEV event classification, both the vanilla untrained BIOT in the fully-supervised setting and the pretrained BIOT model in the fine-tuning setting outperformed all the supervised baselines.
@@
EEGFormer: This model was evaluated on five different scalp EEG datasets, with four fully-supervised models (i.e., EEGNet [74], TCN [144], EEG-GNN [145], and GraphS4mer [146]) and BrainBERT as baselines along with three variations of EEGFormer with different sizes. Fine-tuned EEGFormer outperformed all baselines in all the tasks except one, where its performance was marginally lower than EEG-GNN. Their evaluations also showed that smaller version of EEGFormer demonstrated more stability in some downstream evaluations.
@@
LaBraM: This model was evaluated on four scalp EEG datasets, two clinical (TUAB and TUEV) and two non-clinical (SEED-V and MoBI), while MoBI being the only external evaluation. The evaluations considered several supervised models, SPaRCNet [139], ContraWR [140], CNN-Transformer [141], FFCL [142], ST-Transformer [143], and BIOT as baselines. All variations of LaBraM (Base, Large, and Huge) with fine-tuning outperformed all the baselines in all-4 tasks, where BIOT also performed better than fully-supervised models except in MoBI. The performance gap between LaBraM and BIOT was narrower in the seizure detection task, but wider in the event classification task.
@@
Mentality: This model was evaluated on the TUSZ seizure detection dataset, which was also used for pretraining. The evaluation compared an untrained Mentality model trained in a fully-supervised setting and a pretrained Mentality model linear probed with two linear layers, which showed that pretraining is beneficial in this task.
@@
NeuroLM: The evaluation compared this model fine-tuned jointly on six tasks using a prompt-based tuning with seven baselines fine-tuned individually per task. The evaluations considered supervised models SPaRCNet [139], ContraWR [140], CNN-Transformer [141], FFCL [142], ST-Transformer [143], BIOT, and LaBraM as baselines. Their evaluations showed that joint fine-tuning using prompting is feasible and that provides reasonable performance all tasks. However, their evaluations also showed that NeuroLM did not perform as well as the LaBraM model on any of the tasks, where LaBraM was fine-tuned separately for each task. Additional evaluation using different model sizes of NeuroLM showed that performances on TUAB and TUEV were not sensitive to model size, whereas the performances on HMC and Workload tasks reduced with larger model sizes.
@@
FoME: This model was evaluated on 7 tasks across 4 datasets (MAYO, FNUSA, SEED, SleepEDFx), and evaluations considered several fully-supervised models (LSTM [147], ConvNeXt [148]), general time series models (PatchTST, TimesNet) and some EEG-FMs (BrainBERT, Neuro-GPT, LaBraM) as baselines. Evaluations showed that, in tasks such as emotion classification, sleep staging, and seizure detection, FoME performed competitively or better than all the baselines, whereas, its performance was significantly better in signal forecasting tasks. Interestingly, the evaluations also demonstrated that general time series models, such as PatchTST and TimesNet, perform reasonably well on tasks such as seizure detection with fine-tuning, despite not being pretrained using brain signals.
@@
BrainWave: This model was evaluated using multiple scalp and intracranial EEG datasets and was compared against LaBraM – a scalp-EEG-FM, BrainBERT – an intracranial EEG-FM, and MOMENT[149] – a general time series FM. Evaluations considered different criteria, including class prototype-based few-shot classification and cross-domain transfer learning across subjects, sites, and diagnostic subtypes. Results showed that BrainWave significantly outperformed all the baselines in all settings, with an average improvement of 0.21 AUC points across 12 tasks, establishing its superiority in challenging zero-shot and OOD evaluations.
@@
Evaluations on common tasks: We observed that two datasets, TUAB for abnormal EEG classification (two classes) and TUEV for EEG event detection (six classes), have been widely utilized for downstream evaluation. Note that both TUAB and TUEV are scalp EEG datasets derived from the larger TUH EEG corpus. In Figure 3(a)-3(b), we compare the performances reported on some common clinical tasks. We also analyzed the impact of the pretraining data size in the x-axis. For comparison, we use AUC and F1 scores as performance metrics as they are the most commonly reported metrics for these two tasks. Linear probing results were only available in LaBraM and are indicated with a triangle in the figure. The rest of the results are based on fine-tuning experiments performed by LaBraM, NeuroLM, EEGFormer, and BIOT. We find that AUC values ranged $0.86-0.92$ for TUAB abnormal EEG classification and F1-scores ranged $0.70-0.85$ for TUEV event classification, highlighting a significant performance variation across models. These comparisons also suggested that the size of the pretraining data has a minimal-to-negative relation to model performance on these two tasks.
@@
Figure 3: EEG-FM performances on TUAB and TUEV classifications. The performance of FoME in 3(b) is shown using a line since the pretraining data size was not available. Channel-hours are on a $10^{3}$ scale. All scores represents fine-tuned model performance except for triangular markers that represents linear probed model performance.

### v3

@@ 4.1 Preparation and Representation of Input Data
SD-71 [data:xiang2024resting] (Sleep deprivation detection - binary) $eval^{\#}$
ADHD-Adult [data:trinh2023task] (ADHD classification - binary) $eval^{\#}$
ADHD-Child [data:motie2020eeg] (ADHD classification - binary) $eval^{\#}$
Schizophrenia-28 [data:olejarczyk2017graph] (Schizophrenia classification - binary) $eval^{\#}$
Depression-122 [data:ds003478:1.1.0] (Depression classification - binary) $eval^{\#}$
MDD-64 [data:mumtaz2016mdd] (Major depression detection - binary) $eval^{\#}$
AD-65 [data:miltiadous2023dataset] (Alzheimer’s disease classification - binary) $eval^{\#}$
BCI Competition IV-1 [data:blankertz2007non] (Movement classification - binary) traintrain traintrain
Emobrain [data:savran06_einterface] (Emotion classification - multiclass) traintrain traintrain
Grasp/Lift Challenge [data:luciw2014multi] (Grasp classification/regression) traintrain traintrain
Inria BCI Challenge [data:margaux2012objective] (Next letter prediction - multiclass) traintrain traintrain
EEG Motor Imagery [data:schalk2004bci2000] (Classification - multiclass) traintrain traintrain traintrain
Visual Categorization EEG [data:trujillo2019mental] (Classification - multiclass) traintrain traintrain
Resting-state EEG [data:trujillo2017effect] (Eye state classification - binary) traintrain traintrain
SEED Series [data:zheng2015investigating, data:zheng2018emotionmeter, data:liu2022identifying] (Emotion classification - multiclass) $^{*} train/evaltrain/eval^{*}$ $train/eval^{*}$
SPIS Resting State EEG [data:torkamani2020prediction] (CVS and HRT regression) traintrain traintrain
Brain-Invaders [data:korczowski2019brain] (Target classification - binary) traintrain traintrain
Multiple data sources [data:jiang2021discriminating, data:jiang2023multimodal, data:luo2022multimodal, data:li2021discrimination, data:tao2020emotion] (Unknown tasks) traintrain traintrain
MoBI [data:he2018mobile] (Gait angle regression) $eval^{\#}$
BCI Competition IV Dataset 2a [data:brunner2008bci] $eval^{\#}$
Workload [data:zyma2019electroencephalograms] (Workload classification - binary) $eval^{\#}$
@@
Brain TreeBank [data:wang2024brain] (EEG classification - multiclass) $train/eval^{*}$
@@
Private dataset collection without source (Unknown) traintrain
MAYO [data:nejedly2020multicenter] (Pathology detection & events classification) $eval^{\#}$ $train/eval^{*}$ $eval^{\#}$
FNUSA [data:nejedly2020multicenter] (Pathology detection & events classification) $eval^{\#}$ $train/eval^{*}$ $eval^{\#}$
CCEP [data:ds004080:1.2.4] (Seizure onset zone localization - binary) traintrain
@@ 4.2 Modeling and Pretraining
Patching and context length: The temporal resolution of tokens or patches, i.e., how much EEG data each patch contains, varies between models, reflecting different design choices in model architecture and target tasks. Brant and FoME used patches corresponding to six seconds of EEG, while BrainBERT used 5-second STFT/Superlet windows. Models such as LaBraM, NeuroLM, BIOT, and BrainWave adopted a finer resolution, using 1-second segments per patch, whereas Neuro-GPT utilized a patch size of two seconds. Similarly, context length also varied among EEG-FMs. FoME was trained on 90-second segments, each represented by a sequence of 15 tokens. LaBraM could take up to 256 tokens as input, with the number of temporal tokens depending on the number of channels used. Neuro-GPT utilized a 32-token context window with overlapping segments, covering $\sim$57.8 seconds of EEG. BIOT utilized 19 tokens per channel and EEGFormer utilized 12 seconds of EEG context as input.
Architectural components: The model architecture and training objectives play a central role in determining the effectiveness of a foundation model in extracting information from the input data [bommasani2021opportunities]. EEG-FMs utilized various components at different stages of the pipeline, including tokenizers, local representation learning modules, spatial and/or temporal attention modules to learn dependencies, and task heads. Almost all the EEG-FMs we review in this article are transformer-based, except Mentality, which uses a combination of convolutional layers and a mamba-based architecture. In the transformer-based models, the input signal is divided into a fixed number of small patches/tokens to construct the input sequence. Some models that used raw time-series inputs leveraged convolutional layers to learn morphological features (e.g., Neuro-GPT, LaBraM, Mentality, NeuroLM).
Neural tokenizers: Some models, such as LaBraM, NeuroLM, and EEGFormer, integrated separate tokenizers to transform the input signals into discrete tokens or codes. These EEG-FMs adapted a VQ-VAE [van2017neural]-like architecture to train a neural tokenizer, which mapped input patches into a fixed set of discrete embeddings called the codebook. During tokenizer training, latent features are mapped to nearest-neighbor embeddings from the codebook to reconstruct known features of the input, such as power and phase in LaBraM, and temporal and spectral features in NeuroLM. The models used neural codebooks to reconstruct raw time-series inputs using a transformer-like architecture. While NeuroLM and EEGFormer used codebook embeddings to map encoder outputs to discrete codes prior to decoding, LaBraM used codebook embeddings of masked input patches as targets during model pretraining.
Spatial encoding: EEG-FMs implemented different spatial encoding strategies, although they differ in how explicitly they handle spatial information. Brant and FoME, for example, used transformer encoders with spatial attention but without dedicated spatial positional encodings. In contrast, LaBraM, BIOT, and NeuroLM injected spatial information through explicit spatial encodings to enhance positional encoding and capture the topographic layout of EEG channels. Among them, only BIOT supported multiple electrode configurations, while LaBraM and NeuroLM are limited to a fixed 10–20 layout for spatial encoding. Models such as Neuro-GPT and Mentality learned spatial dependencies implicitly using convolutional layers, without dedicated spatial embeddings. BrainWave incorporated spatial attention mechanisms to model inter-channel dependencies directly. In contrast, BrainBERT and EEGFormer omit spatial modeling altogether, using only transformer layers without convolution or spatial priors.
Self-supervision: SSL approaches are key components that enable the development of foundation models using massive unlabeled datasets [bommasani2021opportunities]. Most EEG-FMs (seven out of ten) employed reconstruction of masked data as the primary SSL approach. EEGFormer model was pretrained by minimizing the reconstruction loss between input signals and signals decoded from codebook mappings of transformer-encoded input patches. NeuroLM employed an autoregressive approach to predict future patches from past patches, including an EEG-text alignment objective, and BIOT adopted a contrastive learning approach where the embeddings of masked inputs and the same inputs with different augmentations are minimized (i.e., the SimCLR [chen2020simple] pretraining approach). Only one model — Neuro-GPT — employed masked data reconstruction in the latent space. Additionally, the LaBraM model employed a masked-signal reconstruction approach; however, the codebook embeddings of masked input patches from a previously trained neural tokenizer were used as reconstruction targets.
Masking: The granularity of the learned representations in the EEG-FMs is influenced by the type and extent of masking used during self-supervised pretraining [he2022masked]. In models based on masked-signal reconstruction, masking is typically applied to random patches across temporal and spatial dimensions, as opposed to masking entire channels or full time segments. For example, Brant and FoME applied a $\sim$40% uniform random masking, zeroing out the masked input signal patches across time and channel dimensions. LaBraM similarly utilized a 50% patch-level masking, while BrainWave, despite operating in the spectrogram domain, performed masking entire spectrograms of several channels. On the other hand, BrainBERT applied a more fine-grained masking strategy, masking 5% from the time domain and 5% from the frequency domain, effectively masking $\sim$10% of the input spectrogram. Mentality applied random channel masking, targeting specific channels rather than patches. NeuroLM introduced a stair-step masking scheme, in which next token is auto-regressively predicted in a channel-conditioned manner, whereas Neuro-GPT utilized a causal masking approach in the latent space, consistent with autoregressive generation.
@@
Model scale: The number of trainable parameters (or weights) varied significantly between the ten EEG-FMs, ranging from 3.3M in BIOT to 1.7B in NeuroLM, although most models fall within the 200-500M parameter range. Some EEG-FMs, such as FoME, LaBraM, and NeuroLM, were developed at multiple scales with the same architecture, allowing users to choose a model scale according to their specific resource constraints.
@@ 4.3 Model Evaluations
Downstream task evaluations: Performance on downstream tasks post-adaptation was the primary metric used for evaluation in all the EEG-FMs. However, the clinical and non-clinical downstream tasks used for evaluation and the adaptation techniques varied significantly between them. The various datasets used for evaluation are shown in Table 2. While most EEG-FM evaluations were restricted to either scalp or intracranial EEG, two models, namely FoME and BrainWave, were evaluated on both modalities. Furthermore, most FMs used common adaptation techniques, namely linear probing and fine-tuning, while some used prototype-based few-shot evaluation and instruction tuning.
@@
Almost all EEG-FMs, except Mentality, investigated fine-tuning for downstream task evaluations, in which the pretext task decoder is discarded but the encoder is trained further using task-specific labels. On the other hand, five EEG-FMs (i.e., Brant, BrainBERT, LaBraM, EEGFormer, and Neuro-GPT) evaluated linear probing, where the encoder is typically frozen, and additional task-specific heads on top of the encoder are trained using labeled downstream data. The BrainWave model was evaluated using a prototype-based few-shot learning approach, where class prototypes were created by averaging the latent representations of a few examples per class, and downstream classification was performed by measuring similarity to those prototypes in the learned latent space. Furthermore, NeuroLM, which was pretrained with an EEG-text alignment objective, was evaluated using a text-prompting approach. The model was fine-tuned (instruction-tuned) using an auto-regressive prediction task utilizing EEG-text combined tokens, where the text tokens contained both class prompts and labels. Subsequently, the model was evaluated on various tasks by providing queries containing EEG tokens and prompts, for which the model predicted the class labels. In the following, we briefly summarize the evaluations performed by each EEG-FM and the results.
@@
BrainBERT: This model was trained and evaluated on the same intracranial EEG dataset, Brain TreeBank. The downstream task focused on predicting the features of the movies that the subjects were watching during the EEG recording. Fine-tuning the pretrained BrainBERT model on this task substantially outperformed fully-supervised linear and multi-layer neural network models. Linear probing also achieved Area Under the Receiver Operating Characteristic Curve (AUC) values similar to those of the fully supervised models.
@@
Neuro-GPT: This model was evaluated on a single scalp EEG dataset that was not used during pretraining, with the downstream task being 4-class motor-imagery classification. The fully-supervised baselines considered in this study include BENDR [kostas2021bendr], SVM [oikonomou2017comparison], EEGNet [lawhern2018eegnet], CTCNN [schirrmeister2017deep], CCNN [amin2019deep], and NG-CRAM [zhang2020motor]. When used in a supervised setting without any pretraining, this model performed similarly to other supervised baselines. However, when the pretrained model was fine-tuned for the downstream task, it provided significant improvements in certain configurations. In contrast, linear probing performed worse than fully-supervised baselines.
@@
Brant: This model was evaluated on three unseen intracranial EEG datasets, including one private and two public datasets (MAYO and FNUSA). The downstream tasks evaluated include signal forecasting, imputation, seizure detection, and pathology detection. For all these tasks, the pretrained model was fine-tuned or linear-probed using task labels. The self-supervised pretrained baselines considered in this study included (RP, TS, CPC) [banville2021uncovering], BENDR [kostas2021bendr], MVTS [potter2022unsupervised], and BrainBERT that are designed for brain signals, and CoST [woo2022cost], TF-C [zhang2022self], PatchTST [nieTimeSeriesWorth2023], TS-TCC [eldele2021time] that are designed for general time series. Fully-supervised baselines without any pretraining included handcrafted features-based methods such as spectral power [zhang2015low], rhythmicity spectrogram [handa2021epileptic], and amplitude-integrated EEG [kadivar2019comparison] along with SEEG-Net [wang2022seeg]. In signal forecasting and imputation tasks, fine-tuned Brant consistently outperformed several time series baselines and its own linear probing version. In the seizure detection and pathology detection tasks, both the fine-tuned and linear probed versions of Brant outperformed other supervised and EEG-FM baselines, including BrainBERT. These observations held true in low-labeled settings as well.
@@
BIOT: This model was evaluated entirely on scalp EEG datasets on various clinical tasks. The supervised baselines considered in this study included SPaRCNet [jing2023development], ContraWR [yang2021self], CNN-Transformer [peh2022transformer], FFCL [li2022motor], ST-Transformer [song2021transformer]. Evaluations showed that the finetuned BIOT encoder outperformed several fully-supervised baselines in CHB-MIT seizure detection and TUAB normal/abnormal classification. In two other tasks, namely IIIC seizure classification and TUEV event classification, both the vanilla untrained BIOT in the fully supervised setting and the pretrained BIOT model in the fine-tuning setting outperformed all supervised baselines.
@@
EEGFormer: The three variants of EEGFormer were evaluated against four fully-supervised models (i.e., EEGNet [lawhern2018eegnet], TCN [bai2018empirical], EEG-GNN [tang2021self], and GraphS4mer [tang2023modeling]) and BrainBERT as baselines, using five different scalp EEG datasets/tasks. Fine-tuned EEGFormer outperformed all baselines in all the tasks except one, where its performance was marginally lower than EEG-GNN. Their evaluations also showed that a smaller version of EEGFormer demonstrated less variability in some downstream evaluations.
@@
LaBraM: This model was evaluated on four scalp EEG datasets: two clinical (TUAB and TUEV) and two non-clinical (SEED-V and MoBI), with MoBI as the only external dataset. The evaluations considered several supervised models, SPaRCNet [jing2023development], ContraWR [yang2021self], CNN-Transformer [peh2022transformer], FFCL [li2022motor], ST-Transformer [song2021transformer], and BIOT as baselines. All variations of LaBraM (base, large, and huge) with fine-tuning outperformed baselines in all tasks.
@@
Mentality: This model was evaluated on the TUSZ seizure detection dataset, which was also used for pretraining. The evaluation compared an untrained Mentality model trained in a fully supervised setting with a pretrained Mentality model linearly probed with two linear layers, showing that pretraining is beneficial for this task.
@@
NeuroLM: This model was fine-tuned jointly on six tasks and was compared against seven baselines fine-tuned individually per task. The evaluations considered supervised models SPaRCNet [jing2023development], ContraWR [yang2021self], CNN-Transformer [peh2022transformer], FFCL [li2022motor], ST-Transformer [song2021transformer], BIOT, and LaBraM as baselines. Evaluations showed that joint fine-tuning with prompting is feasible and achieves reasonable performance across all tasks. However, NeuroLM did not perform as well as the LaBraM model on any of the tasks, where LaBraM was fine-tuned separately for each task. Evaluations showed that the TUAB and TUEV performance of NeuroLM variants was not sensitive to model size, whereas the performance on HMC and Workload tasks decreased with larger variants.
@@
FoME: This model was evaluated on 7 tasks across 4 datasets (MAYO, FNUSA, SEED, SleepEDFx), and evaluations considered several fully-supervised models (LSTM [hochreiter1997long], ConvNeXt [liu2022convnet]), general time series models (PatchTST, TimesNet), and some EEG-FMs (BrainBERT, Neuro-GPT, LaBraM) as baselines. Evaluations showed that, in emotion classification, sleep staging, events classification, and pathology detection tasks, FoME performed competitively or better than all baselines, whereas its performance was significantly better in signal forecasting tasks. Interestingly, the evaluations also demonstrated that general time-series models, such as PatchTST and TimesNet, perform reasonably well on pathology detection after fine-tuning, despite not being pretrained on brain signals.
@@
BrainWave: This model was evaluated using multiple scalp and intracranial EEG datasets and was compared against LaBraM – a scalp-EEG-FM, BrainBERT – an intracranial EEG-FM, and MOMENT [goswami2024moment] – a general time series FM. Evaluations considered different criteria, including class prototype-based few-shot classification and cross-domain transfer learning across subjects, sites, and diagnostic subtypes. Results showed that BrainWave significantly outperformed all the baselines in all settings, with an average improvement of 0.21 AUC points across 12 tasks, establishing its superiority in challenging out-of-distribution evaluations.
@@
Evaluations on common tasks: We observed that two datasets, TUAB for abnormal EEG classification (two classes) and TUEV for EEG event detection (six classes), have been widely utilized for downstream evaluation. Note that both TUAB and TUEV are scalp EEG datasets derived from the larger TUH EEG corpus. In Figure 3(a)-3(b), we compare the performances reported on these common tasks. We also analyzed the impact of the pretraining data size on the x-axis. For comparison, we use AUC and F1 scores as performance metrics, as they are the most commonly reported metrics for these two tasks. Linear probing results were available only in LaBraM and are indicated by a triangle in the figure. The remaining results are based on fine-tuning experiments conducted by LaBraM, NeuroLM, EEGFormer, and BIOT. We find that AUC values ranged 0.860.920.86-0.92 for TUAB abnormal EEG classification and F1-scores ranged 0.700.830.70-0.83 for TUEV event classification, highlighting significant performance variation across models. These comparisons also suggested that increasing the size of the pretraining dataset did not necessarily improve model performance on these two tasks.
@@
Figure 3: Performances on common tasks. Here we compare EEG foundation models based on their performance on the common Temple University EEG Corpus tasks – TUAB (abnormal EEG classification) and TUEV (event classification) – along with the size of the dataset used for pretraining. The performance of FoME in 3(b) is shown as a line because the pretraining data size was unavailable. Channel-hours are on a $10^{3}$ scale. All scores represent fine-tuned model performance, except for the triangular markers, which represent linear-probed model performance.